A pendulum clock giving correct time at a place where g = 9.800 ms-2 is taken to another place where it loses 24 seconds during 24 hours. Find the value of g at this new place.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. . For the pendulum, T_1/T_2 = √(g_2/g_1 ) Given that, T1 = 2sec, g1 = 9.8m/s2 T2 = (24 × 3600)/(((24 × 3600 - 24)/2) ) = 2 × 3600/3599 Now, g^2/g_1 = (T_1/T_2 )^2 ∴ g2 = (9.8) (3599/3600)^2 = 9.795m/s2